# Sharp Thresholds for Temporal Motifs and Doubling Time in Random Temporal Graphs

### v1

@@ 4.1.1 First Moment
With that established, we may turn our attention to construction of the actual bound. The observant reader may notice that the expectations obtained above will tend to zero whenever $\delta(n)=o(n^{-\rho_{H}})$ and $|V_{H}|$ is fixed. In fact Lemma˜4.4 is sufficient to show the lower bounds of Theorem˜3.1 and Theorem˜3.2 via a straightforward application of the first moment method. Unfortunately, the same cannot be said for Theorem˜3.4. Indeed, even a cursory inspection of the problem will reveal that $C^{(\delta)}_{S_{1},\tau_{1}}$ and $C^{(\delta)}_{S_{2},\tau_{2}}$ have a strong positive association whenever $\tau_{1}$ and $\tau_{2}$ share labels or even when $S_{1}$ and $S_{2}$ share a single edge. Consequently, the first moment produces an extremely weak bound that explodes whenever $r>1$ and the size of $H$ is unbounded. In order to control these dependencies and limit the contribution of additional occurrences after the first, we make use of a slightly more sophisticated bound based on the ratio of the expectation and the expectation conditioned on being non-zero.
@@ 4.1.2 Second Moment
Once again we wish to avoid a direct calculation of a lower bound on the probability of a given motif occurring, and so our strategy will be to use the second moment method. Unfortunately, we do not find a unified over-arching bound sufficient to obtain our results for both fixed and growing motifs, as positive association rears its head once more (although this time as an obstacle for the fixed case). In particular, when two potential occurrences of a motif share a footprint they remain dependent even if they do not share any labels, due to the inherent randomness of the label multiplicity distribution. This presents an issue whenever $\psi$ is a non-degenerate non-Bernoulli distribution, as we will retain terms in our summations corresponding to such pairs. These terms contribute an additional $\delta$ factor to our calculations, which can be absorbed in the case of Theorem˜3.4 where $\delta$ is constant but a naive application of the second moment would lead to asymptotically weaker bounds in the case of Theorem˜3.1 and Theorem˜3.2 where $\delta$ is decreasing. In order to avoid this pitfall, we shall make use of two bounds, one for the case where $\psi$ is a Bernoulli random variable and one for the more general but still bounded case. We shall warm up with the considerably easier former case.
@@
Let $\mathcal{G}=(V,E,\lambda)$ be a sample from $\Gamma_{[n]}(\psi)$ for $\psi$ a Bernoulli random variable then for any $0<\delta<1$,
@@
With that established, we can present the following more general result, in which we have to be much more careful to control the dependence caused by the label multiplicity distribution. As a consequence, the following lemma is restricted to the case where $\psi$ is almost surely bounded.
@@ 4.2 Existence Thresholds
The opposing bounds are somewhat more challenging. In both cases, we are only able to obtain a weak variant of our desired result via a naive application of the second moment method. For fixed motifs, we inherit the restriction of $\psi$ to Bernoulli distributions from Lemma˜4.6, and for the largest unordered clique, we similarly find ourselves restricted to almost surely bounded label multiplicity distributions. However, these restricted versions actually turn out to be sufficient for our purposes. The key idea is to apply Lemma˜4.3 and exploit the relationship between samples from multiple instances of our model with different (and stochastically ordered) label multiplicity distributions. On the one hand, in the case of fixed motifs, observe that any non-negative non-degenerate discrete distribution $\psi$ over $\mathbb{N}$ stochastically dominates some Bernoulli distribution $\sigma$. Since we know from Lemma˜4.3 that the number of instances of a temporal motif in two couple samples from the continuous (resp. discrete) model with different label multiplicity distributions inherit the stochastic order of their corresponding label multiplicity distributions, it follows that the probability of a given temporal motif existing in the sample from $\Gamma_{[n]}(\psi)$ is lower bounded by the probability of one existing in the sample from $\Gamma_{[n]}(\sigma)$. Combining this with the bound from the second moment method, even when restricted only to Bernoulli label multiplicity distributions is then sufficient to obtain our threshold. On the other hand, we have to be a bit more careful in the case of the largest unordered clique, as the error introduced from the discrepancy between the actual label multiplicity distribution and a Bernoulli lower bound is too high. Fortunately, for any valid choice $\psi$ of label multiplicity distribution, we can find a second almost surely bounded distribution $\chi$ such that $\psi$ stochastically dominates $\chi$ and $\chi$ has first and second moments arbitrarily close to those of $\psi$. In particular, for a suitable choice of $\chi$, we are able to achieve the same bound in Theorem˜3.4 for both samples from $\Gamma_{n}(\psi)$ and $\Gamma_{[n]}(\chi)$. The result then follows from the same stochastic order argument as for fixed motifs.
@@ 5 Doubling Times
For a temporal graph $\mathcal{G}=(V,E,\lambda)$ where no two edges share the same label, we define $\mathcal{G}^{O}=(V,E,\lambda^{\prime})$ the order discretization of $\mathcal{G}$ such that $\lambda^{\prime}(e)$ corresponds to the set of relative positions of each label in $\lambda(e)$ in the overall set of labels. Furthermore, for $0\leq a<b$ we define $\mathcal{G}^{O}_{[a,b]}=(V,E_{[a,b]},\lambda^{\prime}_{[a,b]})$ to be $\mathcal{G}^{O}$ with all labels outside of $[a,b]$ removed.
Since the number of labels falling within a given interval of a sample from $\Gamma_{[n]}(\psi)$ is highly concentrated, we are easily able to convert between times in a sample from $\Gamma_{[n](\psi}$ and its order discretization (see Lemma˜9.7).
We will begin with our argument for obtaining an upper bound on the doubling time. However, before we begin we will need a few preliminary ideas. A key property of our model is that it is in some sense “time reversible” as the probability of any set of edge labels appearing in a given order is the same as that of them falling in the reverse (or in fact any) order. This provides us with a very useful tool as we are able to consider our normal temporal graph and variants of it where we have partially reversed its labeling interchangeably. The next definition and lemma (see also Definition˜2.6) exploits this property to reduce bounding the doubling time of small sets of vertices to bounding the doubling time of large sets of vertices in the partially reversed graph.
@@
  • $\textbf{Double}^{-}(\mathcal{G})=\max_{S\subset V:|S|\leq\frac{n}{4}}\left(\text{Double}_{\mathcal{G}}(S)\right)$ to be doubling time over small sets.
  • $\textbf{Double}^{+}(\mathcal{G})=\max_{S\subset V:|S|>\frac{n}{4}}\left(\text{Double}_{\mathcal{G}}(S)\right)$ to be the doubling time over large sets.
@@
While stated for arbitrary temporal graphs, in our setting, given the equal probability of sampling both $\mathcal{G}$ and $(\mathcal{R}(\mathcal{G},\textbf{Double}^{+}(\mathcal{G}))$, this turns any upper bound on the doubling time of large sets into an upper bound on the doubling time of small sets. Our second key preliminary is an upper bound on the time taken for the reachability ball centered on any specific set of vertices to grow to a certain a size, obtained from a carefully constructed coupling between the total time between expansions of the reachability ball and a sum of geometric random variables. We defer its precise (and somewhat involved) statement to the detailed exposition (see Lemma˜9.3).
@@
The proof of this result consists largely of applying Lemma˜9.3, along with a concentration inequality and Lemma˜5.3 in order to deal with the case of small sets. More specifically, observe that once all vertices can be reached by at least $\frac{3n}{4}$ other vertices, every set of size at least $\frac{n}{4}$ can reach all vertices. Thus, once this has occurred the reachability ball of all large sets of vertices must have doubled in size. In order to bound this time, we first show via an application of Lemma˜9.3 and Markov’s inequality that almost all vertices can be reached by $\frac{3n}{4}$ other vertices by roughly time $n\log{n}$. From a second application of Lemma 9.3 we find that the few remaining vertices are reached from the already well connected vertices within only a further approximately $n\log{n}$ rounds. The more general result then follows from using Lemma˜5.3 to extend the bound to cover all smaller sets of vertices.
Finally, we turn our attention to deriving a lower bound to match the upper bound of Lemma˜5.5. However, this time, our life will be significantly easier, as it suffices to consider only the doubling time of sets containing precisely $\frac{n}{2}$ vertices. Analogously to the upper bound, we first establish a lower bound on the doubling time of any specific set of $\frac{n}{2}$ vertices by coupling with a sum of geometric random variables (see Lemma˜9.5). We can then apply this to prove a lower bound on the doubling time of $\mathcal{G}^{O}$.
@@ 6 Related Work
Most research concerning temporal motifs has focused on either algorithmic techniques for counting or estimating the number of temporal motifs, or the application of these methodologies for the analysis of real networks. A frequent issue in the latter is the construction of a “null” model to compare the relative frequency of motifs against. Despite this there has been very limited research into the distribution of motifs in random temporal graphs. A notable exception to this is [AnalyticalModels], which derives algorithms for calculating the expectation and the variance of the number of motifs appearing in a proposed random temporal graph model. While closely related, their work differs from ours in several ways. Firstly, they focus primarily on the construction of a model that can be statistically fit to data, and so use a stochastic block model footprint graph with Poisson distributed inter edge appearance times. This permits them a greater degree of topological richness than is obtained in our simpler model, however the distribution of the labels on each edge is more restricted. Secondly, they focus primarily on providing an algorithm for the calculation of the first and second moment, rather than their application to find existence thresholds. Thirdly, they work only in the continuous setting and do not consider anything comparable to our discrete model. An interesting open problem, would be closing this gap, in particular providing equivalent results in a unified model capturing both the topological expressivity of the stochastic block model and temporal expressivity of our label multiplicity distribution based approach. While not presented as a result on temporal motifs, [Cliques] investigates the growth of the largest $\delta$-temporal clique, which implicitly provides thresholds for motifs with a clique footprint and a trivial partial order. In this work, we strictly generalize these results to the more powerful random graph models we consider.
@@
Until recently, as discussed in the excellent survey of [SharpThresholds], reachability problems in random temporal graphs have primarily been studied in other fields, such as population protocols, the gossip model or the random edge ordering model. For example, in population protocols the reachability properties of the underlying temporal graph underpin two of the model’s most fundamental lower bounds on the time complexity of broadcast [DBLP:conf/nca/MocquardSRA16] and leader election [DBLP:journals/corr/abs-1906-11121]. However, recently, in part due to the seminal work of [SharpThresholds], there has been growing interest in the reachability properties of random temporal graphs from both a structural and algorithmic perspective. While other models have been considered (see for example [baguley2026temporal, https://doi.org/10.1002/rsa.70040, brandenberger2025temporal]), there has been a focus on establishing the properties of the so-called “Random Simple Temporal Graph” defined as a binomial random graph with a linear order over its edges. Sharp thresholds for have since been established at: $\frac{\log{n}}{n}$ for point to point to point reachability [SharpThresholds] and the existence of a giant [becker2022giant] (and even super constant [atamanchuk2024size]) connected-component; $\frac{2\log{n}}{n}$ for the existence of temporal sources and sinks; $\frac{3\log{n}}{n}$ for full temporal connectivity[SharpThresholds]; and $\frac{4\log{n}}{n}$ for the existence of pivotal (and non-sharp optimal) spanners [SharpThresholds]. Additionally, there have been several results concerning various characteristics of the length of time-respecting paths in the RSTG and their connections to connectivity [angel2020long, broutin2024increasing]. Our treatment of doubling time shares the most similarity with [becker2022giant] as similarly to them we require bounds on the growth of reachability forests, however rather than using a foremost forest (as in [becker2022giant, SharpThresholds]) or the random recursive tree (as in [broutin2024increasing]), we instead derive the times directly via a simple coupling with a sum of geometric random variables. An additional distinction is our choice of models which both have asymptotically constant density and are not necessarily simple. Despite this, when restricted to the interval $[a,a+p]$ with a single label per edge, sampling from our continuous model is equivalent to sampling from the RSTG with parameter $p$.
@@ 7 Auxiliary Results
We will need some of the following results.
@@
Let $X$ be a non-negative random variable. Then for any $\delta>0$,
@@
Let $(X_{t})_{t\in\mathbb{N}_{0}}$, be a stochastic process, adapted to a filtration $(\mathcal{F}_{t})_{t\in\mathbb{N}_{0}}$, over some state space $S\subseteq\mathbb{R}$. For some $a\geq 0$, such that $\{x\in S|x\leq a\}$ is absorbing, let $T_{a}=\min\{t\geq 0:X_{t}\leq a\}$. Moreover, let $g:S\to\mathbb{R}_{\geq 0}$ be a function such that $g(0)=0$ and $g$ is monotonically increasing on $\mathbb{R}_{\geq 0}$. If there exists $\lambda>0$ and a value $\beta>0$ such that the following holds for all $t\in\mathbb{N}_{0}$,
@@
Let $X$ be a random variable with a continuous cumulative distribution function supported entirely on $(0,1]$, such that $E[X|X>0]$ and $E[X^{2}|X>0]$ both exist. Further let $Y,Z\sim\text{Ber}(X)$ be independent samples from a Bernoulli random variable parameterized by the same sample from $X$, then
@@
We have the following for $p(x)$ the probability density function of $X|X>0$,

### v2

@@ 4.1.1 First Moment
With that established, we may turn our attention to construction of the actual bound. The observant reader may notice that the expectations obtained above will tend to zero whenever $\delta(n)=o(n^{-\rho_{H}})$ and $|V_{H}|$ is fixed. In fact, Lemma˜4.4 is sufficient to show the lower bounds of Theorem˜3.1 and Theorem˜3.2 via a straightforward application of the first moment method. Unfortunately, the same cannot be said for Theorem˜3.4. Indeed, even a cursory inspection of the problem will reveal that $C^{(\delta)}_{S_{1},\tau_{1}}$ and $C^{(\delta)}_{S_{2},\tau_{2}}$ have a strong positive association whenever $\tau_{1}$ and $\tau_{2}$ share labels or even when $S_{1}$ and $S_{2}$ share a single edge. Consequently, the first moment produces an extremely weak bound that explodes whenever $r>1$ and the size of $H$ is unbounded. In order to control these dependencies and limit the contribution of additional occurrences after the first, we make use of a slightly more sophisticated bound based on the ratio of the expectation and the expectation conditioned on being non-zero.
@@ 4.1.2 Second Moment
Once again we wish to avoid a direct calculation of a lower bound on the probability of a given motif occurring, and so our strategy will be to use the second moment method. Unfortunately, we do not find a unified overarching bound sufficient to obtain our results for both fixed and growing motifs, as positive association rears its head once more (although this time as an obstacle for the fixed case). In particular, when two potential occurrences of a motif share a footprint they remain dependent even if they do not share any labels, due to the inherent randomness of the label multiplicity distribution. This presents an issue whenever $\psi$ is a non-degenerate non-Bernoulli distribution, as we will retain terms in our summations corresponding to such pairs. These terms contribute an additional $\delta$ factor to our calculations, which can be absorbed in the case of Theorem˜3.4 where $\delta$ is constant but a naive application of the second moment would lead to asymptotically weaker bounds in the case of Theorem˜3.1 and Theorem˜3.2 where $\delta$ is decreasing. In order to avoid this pitfall, we shall make use of two bounds, one for the case where $\psi$ is a Bernoulli random variable and one for the more general but still bounded case. We shall warm up with the considerably easier former case.
@@
Let $\mathcal{G}=(V,E,\lambda)$ be a sample from $\Gamma_{[n]}(\psi)$, for $\psi$ a Bernoulli random variable, then for any $0<\delta<1$,
@@
With that established, we can present the following more general result, in which we have to be much more careful to control the dependence caused by the label multiplicity distribution. Consequently, the following lemma is restricted to the case where $\psi$ is almost surely bounded.
@@ 4.2 Existence Thresholds
The opposing bounds are somewhat more challenging. In both cases, we are only able to obtain a weak variant of our desired result via a naive application of the second moment method. For fixed motifs, we inherit the restriction of $\psi$ to Bernoulli distributions from Lemma˜4.6, and for the largest unordered clique, we similarly find ourselves restricted to almost surely bounded label multiplicity distributions. However, these restricted versions actually turn out to be sufficient for our purposes. The key idea is to apply Lemma˜4.3 and exploit the relationship between samples from multiple instances of our model with different (and stochastically ordered) label multiplicity distributions. On the one hand, in the case of fixed motifs, observe that any non-negative non-degenerate discrete distribution $\psi$ over $\mathbb{N}$ stochastically dominates some Bernoulli distribution $\sigma$. Since we know from Lemma˜4.3 that the number of instances of a temporal motif in two coupled samples from the continuous (resp. discrete) model with different label multiplicity distributions inherit the stochastic order of their corresponding label multiplicity distributions, it follows that the probability of a given temporal motif existing in the sample from $\Gamma_{[n]}(\psi)$ is lower bounded by the probability of one existing in the sample from $\Gamma_{[n]}(\sigma)$. Combining this with the bound from the second moment method, even when restricted only to Bernoulli label multiplicity distributions, is then sufficient to obtain our threshold. On the other hand, we have to be a bit more careful in the case of the largest unordered clique, as the error introduced from the discrepancy between the actual label multiplicity distribution and a Bernoulli lower bound is too high. Fortunately, for any valid choice $\psi$ of label multiplicity distribution, we can find a second almost surely bounded distribution $\chi$ such that $\psi$ stochastically dominates $\chi$ and $\chi$ has first and second moments arbitrarily close to those of $\psi$. In particular, for a suitable choice of $\chi$, we are able to achieve the same bound in Theorem˜3.4 for both samples from $\Gamma_{n}(\psi)$ and $\Gamma_{[n]}(\chi)$. The result then follows from the same stochastic order argument as for fixed motifs.
@@ 5 Doubling Times
For a temporal graph $\mathcal{G}=(V,E,\lambda)$ where no two edges share the same label, we define $\mathcal{G}^{O}=(V,E,\lambda^{\prime})$ the order discretization of $\mathcal{G}$ such that $\lambda^{O}(e)$ corresponds to the set of relative positions of each label in $\lambda(e)$ in the overall set of labels. Furthermore, for $0\leq a<b$ we define $\mathcal{G}^{O}_{[a,b]}=(V,E_{[a,b]},\lambda^{O}_{[a,b]})$ to be $\mathcal{G}^{O}$ with all labels outside of $[a,b]$ removed.
Since the number of labels falling within a given interval of a sample from $\Gamma_{[n]}(\psi)$ is highly concentrated, we are easily able to convert between times in a sample from $\Gamma_{[n]}(\psi)$ and its order discretization (see Lemma˜9.8).
We will begin with our argument for obtaining an upper bound on the doubling time. However, first, we will need a few preliminary ideas. A key property of our model is that it is in some sense “time reversible” as the probability of any set of edge labels appearing in a given order is the same as that of them falling in the reverse (or in fact any) order. This provides us with a very useful tool as we are able to consider our normal temporal graph and variants of it where we have partially reversed its labeling interchangeably. The next definition and lemma (see also Definition˜2.6) exploits this property to reduce bounding the doubling time of small sets of vertices to bounding the doubling time of large sets of vertices in the partially reversed graph.
@@
  • $\textbf{Double}^{-}(\mathcal{G})=\max_{S\subset V:|S|\leq\frac{n}{4}}\left(\text{Double}_{\mathcal{G}}(S)\right)$ to be the doubling time over small sets.
  • $\textbf{Double}^{+}(\mathcal{G})=\max_{S\subset V:\frac{n}{4}<|S|\leq\frac{n}{2}}\left(\text{Double}_{\mathcal{G}}(S)\right)$ to be the doubling time over large sets.
@@
While stated for arbitrary temporal graphs, in our setting, given the equal probability of sampling both $\mathcal{G}$ and $\mathcal{R}(\mathcal{G},\textbf{Double}^{+}(\mathcal{G}))$, this turns any upper bound on the doubling time of large sets into an upper bound on the doubling time of small sets. Our second key preliminary is an upper bound on the time taken for the reachability ball centered on any specific set of vertices to grow to a certain size, obtained from a carefully constructed coupling between the total time between expansions of the reachability ball and a sum of geometric random variables. We defer its precise (and somewhat involved) statement to the detailed exposition (see Lemma˜9.3).
@@
The proof of this result consists largely of applying Lemma˜9.3, along with a concentration inequality and Lemma˜5.3 in order to deal with the case of small sets. More specifically, observe that once all vertices can be reached by at least $\frac{3n}{4}$ other vertices, every set of size at least $\frac{n}{4}$ can reach all vertices. Thus, once this has occurred the reachability ball of all large sets of vertices must have doubled in size. In order to bound this time, we first show via an application of Lemma˜9.3 and Markov’s inequality that almost all vertices can be reached by $\frac{3n}{4}$ other vertices by roughly time $n\log{n}$. From a second application of Lemma˜9.3 we find that the few remaining vertices are reached from the already well connected vertices within only approximately a further $n\log{n}$ rounds. The more general result then follows from using Lemma˜5.3 to extend the bound to cover all smaller sets of vertices.
Finally, we turn our attention to deriving a lower bound to match the upper bound of Lemma˜5.5. However, this time, our lives will be significantly easier, as it suffices to consider only the doubling time of sets containing precisely $\frac{n}{2}$ vertices. Analogously to the upper bound, we first establish a lower bound on the doubling time of any specific set of $\frac{n}{2}$ vertices by coupling with a sum of geometric random variables (see Lemma˜9.5). We can then apply this to prove a lower bound on the doubling time of $\mathcal{G}^{O}$.
@@ 6 Related Work
Most research concerning temporal motifs has focused on either algorithmic techniques for counting or estimating the number of temporal motifs, or the application of these methodologies for the analysis of real networks [sariyuce2025powerful]. A frequent issue in the latter is the construction of a “null” or “reference” model to compare the relative frequency of motifs against [sariyuce2025powerful, kovanen2011temporal, doi:10.1137/19M1242252]. Despite this there has been very limited research into the distribution of motifs in random temporal graphs. A notable exception to this is [AnalyticalModels], which derives algorithms for calculating the expectation and the variance of the number of motifs appearing in a proposed random temporal graph model. While closely related, their work differs from ours in several ways. Firstly, they focus primarily on the construction of a model that can be statistically fit to data, and so use a stochastic block model footprint graph with Poisson distributed inter-edge appearance times. This permits them a greater degree of topological richness than is obtained in our simpler model, however the distribution of the labels on each edge is more restricted. Secondly, they focus primarily on providing an algorithm for the calculation of the first and second moment, rather than their application to find existence thresholds. Thirdly, they work only in the continuous setting and do not consider anything comparable to our discrete model. An interesting open problem is closing this gap, in particular providing equivalent results in a unified model capturing both the topological expressivity of the stochastic block model and temporal expressivity of our label multiplicity distribution based approach. While not presented as a result on temporal motifs, [Cliques] investigates the growth of the largest $\delta$-temporal clique, which implicitly provides thresholds for motifs with a clique footprint and a trivial partial order. In this work, we strictly generalize these results to the more powerful random graph models we consider.
@@
Until recently, as discussed in [SharpThresholds], reachability problems in random temporal graphs have primarily been studied in other fields, such as population protocols, the gossip model or the random edge ordering model. For example, in population protocols the reachability properties of the underlying temporal graph underpin two of the model’s most fundamental lower bounds: the time complexity of broadcast [DBLP:conf/nca/MocquardSRA16] and leader election [DBLP:journals/corr/abs-1906-11121]. However, recently, in part due to the seminal work of [SharpThresholds], there has been growing interest in the reachability properties of random temporal graphs from both a structural and algorithmic perspective. While other models have been considered (see for example [baguley2026temporal, https://doi.org/10.1002/rsa.70040, brandenberger2025temporal]), there has been a focus on establishing the properties of the so-called “Random Simple Temporal Graph” defined as an Erdős-Rényi random graph with a linear order over its edges. Sharp thresholds have since been established at: $\frac{\log{n}}{n}$ for point-to-point reachability [SharpThresholds] and the existence of a giant [becker2022giant] (or even super constant [atamanchuk2024size]) connected-component; $\frac{2\log{n}}{n}$ for the existence of temporal sources and sinks; $\frac{3\log{n}}{n}$ for full temporal connectivity[SharpThresholds]; and $\frac{4\log{n}}{n}$ for the existence of pivotal (and non-sharply optimal) spanners [SharpThresholds]. Additionally, there have been several results concerning various characteristics of the length of time-respecting paths in the RSTG and their connections to connectivity [angel2020long, broutin2024increasing]. Our treatment of doubling time shares the most similarity with [becker2022giant] as, similarly to there, we require bounds on the growth of reachability forests, however rather than using a foremost forest (as in [becker2022giant, SharpThresholds]) or the random recursive tree (as in [broutin2024increasing]), we instead derive the times directly via a simple coupling with a sum of geometric random variables. An additional distinction is our choice of models which are not necessarily simple and have asymptotically constant density. Despite this, when restricted to the interval $[a,a+p]$ with a single label per edge, sampling from our continuous model is equivalent to sampling from the RSTG with parameter $p$.
@@ 7 Auxiliary Results
Before we can begin with any of the proofs, we will need some of the following well known results.
@@
Let $X$ be a non-negative random variable. Then for any $\delta>1$,
@@
Let $(X_{t})_{t\in\mathbb{N}_{0}}$, be a stochastic process, adapted to a filtration $(\mathcal{F}_{t})_{t\in\mathbb{N}_{0}}$, over some state space $S\subseteq\mathbb{R}_{\geq}0$. For some $a\geq 0$, such that $\{x\in S|x\leq a\}$ is absorbing, let $T_{a}=\min\{t\geq 0:X_{t}\leq a\}$. Moreover, let $g:S\to\mathbb{R}_{\geq 0}$ be a function such that $g(0)=0$ and $g$ is increasing on $\mathbb{R}_{\geq 0}$. If there exists $\lambda>0$ and a value $\beta>0$ such that the following holds for all $t\in\mathbb{N}_{0}$,
@@
Let $X$ be a random variable with a continuous cumulative distribution function supported entirely on $[0,1]$, such that $E[X|X>0]$ and $E[X^{2}|X>0]$ both exist. Further let $Y,Z\sim\text{Ber}(X)$ be independent samples from a Bernoulli random variable parameterized by the same sample from $X$, then
@@
We have the following,